OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -2 A) x = 0 B) x = 2 C) There is no x value as there is no solution. D) x can be any value as there is an infinite number of solutions.

OpenStudy (anonymous):

@team_cuban

OpenStudy (anonymous):

I am sorry I forgot how to do system of equations by substitution unfortunately.

OpenStudy (anonymous):

do you know anyone who is good at these and could help me out?

OpenStudy (anonymous):

@pooja195 @welshfella

OpenStudy (anonymous):

These people may be able to help.

pooja195 (pooja195):

Whats the opposite of 2?

pooja195 (pooja195):

any ideas?

OpenStudy (anonymous):

-2

pooja195 (pooja195):

good so we multiply -2(2x + y = 1) 4x + 2y = -2 \[\huge~\rm~-2(2x + y = 1) \]

pooja195 (pooja195):

Distribute

OpenStudy (anonymous):

I have no idea

pooja195 (pooja195):

-2 x 2 = ?

OpenStudy (anonymous):

-4

pooja195 (pooja195):

-2 x y = ?

OpenStudy (anonymous):

I don't know

pooja195 (pooja195):

-2(2x + y = 1) 4x + 2y = -2 \[\huge~\rm~-2(2x + y = 1) \] \(\huge\color{red}{-4x-2y=-2}\)

pooja195 (pooja195):

−4x−2y=−2 4x+2y=−2 Heres the fun part :D -4+4=? -2+2=? -2+-2=?

OpenStudy (anonymous):

can you do that for me... im still a little lost..

pooja195 (pooja195):

-4+4=?

OpenStudy (anonymous):

0

pooja195 (pooja195):

-2+2=?

OpenStudy (anonymous):

0

pooja195 (pooja195):

-2+-2?

OpenStudy (anonymous):

-4

pooja195 (pooja195):

When we have a number = a DIFFERENT number we have no solution its not possible to solve 0=-4 would make it so that there is no solution

OpenStudy (anonymous):

so C?

pooja195 (pooja195):

ye

OpenStudy (anonymous):

See @pooja195 is amazing in Algebra. And I am great in geometry. GOOD WORK @pooja195

pooja195 (pooja195):

:) thanks :)

OpenStudy (anonymous):

No Problem